imqwerty (imqwerty):

Little fun question! :D Find the maximum value of \(\Large y= \frac{x^4-x^2}{x^6+2x^3-1} \) for x>1

OpenStudy (ebayminer126):

?

OpenStudy (faiqraees):

derivatives?

OpenStudy (ebayminer126):

Algebra sucks. .-.

imqwerty (imqwerty):

taking the derivative will be too long :)

imqwerty (imqwerty):

Find the maximum value of \(\Large y= \frac{x^4-x^2}{x^6+2x^3-1}\) for x>1

OpenStudy (mayankdevnani):

i think max value would be 1/6 ! m not sure

imqwerty (imqwerty):

method is important :) how u got to 1/6 ?

OpenStudy (mayankdevnani):

is 1/6 in options ?

imqwerty (imqwerty):

subjective lol

OpenStudy (mayankdevnani):

oh!

imqwerty (imqwerty):

okay these are your options- a)1/6 b)3/4 c)2/5 d)1/3

OpenStudy (mayankdevnani):

option A

imqwerty (imqwerty):

how?

OpenStudy (mayankdevnani):

i used LIMIT CONCEPT here

OpenStudy (mayankdevnani):

first i check lim tending to x-> 1-

OpenStudy (mayankdevnani):

\[\large \bf \lim_{x \rightarrow 1^{1-}}\frac{x^4-x^2}{x^6+2x^3-1}=- \infty \]

OpenStudy (mayankdevnani):

but we don't want this ! because we want for x>1

OpenStudy (mayankdevnani):

so i checked for :- \[\large \bf \lim_{x \rightarrow 1^{+}}\frac{x^4-x^2}{x^6+2x^3-1}=\frac{1}{6}\]

OpenStudy (mayankdevnani):

but i want to check if this is maxima or not

OpenStudy (mayankdevnani):

so i put x=2, i get value less than 1/6

OpenStudy (mayankdevnani):

so this is local maxima

OpenStudy (mayankdevnani):

so 1/6 is our final answer

OpenStudy (agent0smith):

@mayankdevnani I'm curious, how did you get 1/6 from the limit approaching x=1 from the right?

OpenStudy (mayankdevnani):

\[\large \bf \frac{0}{0}~form\]

OpenStudy (mayankdevnani):

then apply L HOSPITAL rule

OpenStudy (agent0smith):

Elaborate. The limit exists and is zero...

OpenStudy (agent0smith):

The function has a root at x=1.

OpenStudy (agent0smith):

How'd you get 0/0 form?

OpenStudy (mayankdevnani):

Oh! Gosh i made a super mistake

OpenStudy (mayankdevnani):

@Empty and @ganeshie8

OpenStudy (mayankdevnani):

@whpalmer4

OpenStudy (faiqraees):

1/6 is correct. Done through graphical method

OpenStudy (faiqraees):

@mayankdevnani

OpenStudy (mayankdevnani):

so how did you draw graph ?

OpenStudy (faiqraees):

used calculator to make a table and then observed the change

OpenStudy (mayankdevnani):

oh! but is bad way to solve it

OpenStudy (faiqraees):

solution is solution

OpenStudy (mayankdevnani):

hey,calc is not allowed in examination

OpenStudy (faiqraees):

in mine it is allowed

OpenStudy (mayankdevnani):

oh! ok

OpenStudy (faiqraees):

although separate marks are given for method

OpenStudy (mayankdevnani):

hmm

OpenStudy (empty):

You can factor it, I dunno if that's helpful to you. \[\frac{x^2 (x+1)(x-1)}{(x-1)^2(x^2+x-1)^2}\] Then cancel a term: \[\frac{x^2 (x+1)}{(x-1)(x^2+x-1)^2}\]

OpenStudy (mayankdevnani):

@Empty how this will help us ?

OpenStudy (mayankdevnani):

@ikram002p

OpenStudy (empty):

I just said I dunno if that's helpful lol.

OpenStudy (mayankdevnani):

lol. it doesn't really help me

OpenStudy (empty):

Sorry to hear that lol

OpenStudy (anonymous):

That denominator is part of a modified version of the generating function for the Pell numbers \(P_n\): \[\sum_{n\ge0}P_nt^n=-\frac{t}{t^2+2t-1}\]and setting \(t=x^3\) we get something resembling the given function, \[\sum_{n\ge0}P_nx^{3n}=-\frac{x^3}{x^6+2x^3-1}\] How does this help? Not sure yet, if it even does... However, we can write \[y=\left(x-\frac{1}{x}\right)\mathcal{G}(P_n;x^3)\]where \(\mathcal{G}(a_n;t)\) is the generating function of the sequence \(a_n\) with respect to \(t\).

OpenStudy (agent0smith):

I disagree with this being a "Little fun question!" I thought there was going to be something particularly clever/simple about how to solve it.

OpenStudy (mayankdevnani):

@ikram002p

OpenStudy (ikram002p):

x is real ?

OpenStudy (mayankdevnani):

yep

OpenStudy (ikram002p):

why can't we just differentiate and take root of y' ? so we would got local values, i also think this function is not bounded so that what u mean of "max" is i can't guess if it's the same as maximizing the function. anyway here is wolfram calculation i'm too lazy to do them :P https://www.wolframalpha.com/input/?i=maximize+(+%7Bx%5E4-x%5E2%7D%5C%7Bx%5E6%2B2x%5E3-1%7D)

OpenStudy (mayankdevnani):

but differentiation is too lenghty

OpenStudy (mayankdevnani):

lengthy

OpenStudy (ikram002p):

no it's simple enough to do polynomial over polynomial

OpenStudy (mayankdevnani):

so can you do for me

OpenStudy (ikram002p):

do what ?

OpenStudy (mayankdevnani):

differentiation

OpenStudy (ikram002p):

oh no that is boring :P

imqwerty (imqwerty):

There is another nice solution :)

OpenStudy (mathmath333):

\(\color{white}{\text{math.stackexchange.com/questions/1380427/maximum-value-of-expression}}\)

OpenStudy (agent0smith):

That was not worth waiting 10 days for. @mathmath333 why did you hide the link? It's not like anyone here was still trying to solve this. I'll make it easier, for anyone who's interested: http://math.stackexchange.com/questions/1380427/maximum-value-of-expression

imqwerty (imqwerty):

that was exactly my method^ sry i forgot that to post solution..

OpenStudy (mathmath333):

i thought that people r still trying to solve not to spoil them , it was not hidden as u did find that